# Impact of 3D macro physics and nuclear physics on the $p$ nuclei in $\mathrm{O}$-$\mathrm{C}$ shell mergers

### v1

@@ Abstract
Massive stars (732) — Oxygen burning (1193) — Stellar convective shells (300) — P-process (1195) — Nuclear astrophysics (1129)
journal: ApJsoftware: Data will be made available on https://astrohub.uvic.ca.
1 Introduction
The origin of the solar pattern of the 35 stable $p$ nuclei is a long-standing problem for understanding nucleosynthesis. Burbidge et al. (1957) first identified these isotopes as a distinct group, produced primarily not by the $s$ or $r$ process, but instead by a $p$ process through $(p,\gamma)$, $(\gamma,p)$, and $(\gamma,n)$ reactions occurring during Type II supernovae and possibly Type I supernovae. Based on the first generations of stellar computational models, Arnould (1976) found that the $p$ process could be driven by all photo-disintegration reactions $(\gamma,n)$, $(\gamma,\alpha)$ and $(\gamma,p)$ during the most advanced evolutionary stages of massive stars. Woosley & Howard (1978) found instead that the same photo-disintegrations could build a distribution of $p$ process nuclei during the passage of the supernova shock over the internal progenitor structure, which they called the $\gamma$ process where the production of $p$-process nuclei where they are driven by photo-disintegrations. Following works better defined $\gamma$ process production in Type II supernovae and more in general in core-collapse supernovae (CCSNe e.g., Prantzos et al., 1990; Rayet et al., 1995; Travaglio et al., 2018; Choplin et al., 2022; Roberti et al., 2023, 2024).
@@ 1 Introduction
The $\gamma$ process describes the flow of $(\gamma,n)$, $(\gamma,p)$, and $(\gamma,\alpha)$ reactions on the stable isotopic seeds that are already present in the shell at temperatures of $1.5\leq T\leq 3\times 10^{9}\;\mathrm{K}$ (Rauscher et al., 2013). However, the $\gamma$ process in massive stars underproduces not only the $\mathrm{Mo}$ and $\mathrm{Ru}$, but all $p$ nuclei with $A=90{-}130$ (Arnould & Goriely, 2003; Woosley & Heger, 2007). Travaglio et al. (2011) showed that the $\gamma$ process could produce all $p$ nuclei during a Type Ia supernova from the $s$-process material synthesized during stellar evolution without the underproduction of $\mathrm{{}^{92,94}Mo}$ and $\mathrm{{}^{96,98}Ru}$. Furthermore, Travaglio et al. (2015) found that modifying the distribution of $s$-process material significantly influenced the production of the $p$ nuclei, especially the heaviest ones, and that the lightest three were strongly dependent on the metallicity. Battino et al. (2020) additionally found that the H-flashes of rapidly accreting white dwarfs which undergo the $i$ process could modify the seed distribution to produce $p$ nuclei with $19696<A<$ 196 by the $\gamma$ process during the subsequent SNIa.
Since the first isotopic classification made by Burbidge et al. (1957), it has been found that not all $p$ nuclei in the solar pattern are produced by a single process. Bisterzo et al. (2011) state that $\mathrm{{}^{152}Gd}$, $\mathrm{{}^{164}Er}$, and $\mathrm{{}^{180}Ta}$ have significant contributions of $70.5\%$, $75.5\%$, and $74.5\%$ from the $s$ process. Dillmann et al. (2008) found that $\mathrm{{}^{113}In}$ and $\mathrm{{}^{115}Sn}$ are made by $\beta$-decays after the $r$ process through isomeric states. Goriely et al. (2001) argue that $(\gamma,n)$ was too weak to produce $\mathrm{{}^{138}La}$, and instead that it is made by $\nu_{e}$-capture on $\mathrm{{}^{138}Ba}$ during the CCSN, and Arnould & Goriely (2003) similarly say that $\mathrm{{}^{180m}Ta}$ could also have $\nu$-induced contributions. Sieverding et al. (2018) also found that the $\nu$-process is important for the nucleosynthesis of $\mathrm{{}^{113}In}$, $\mathrm{{}^{138}La}$, and $\mathrm{{}^{180}Ta}$.
The $\mathrm{O}$-shell where $p$ nuclei are produced during a merger is a convective-reactive environment where mixing and nuclear burning timescales are equal (Ritter et al., 2018a; Yadav et al., 2020). If there is significant energy released the flow can be modified (Dimotakis, 2005), such as $\mathrm{H}$-ingestion into $\mathrm{He}$-burning shell (Herwig et al., 2011, 2014) or $\mathrm{O}$-$\mathrm{C}$ shell mergers causing violent mixing (Andrassy et al., 2020; Yadav et al., 2020). The Damköhler number (Dimotakis, 2005) quantifies the ratio of these timescales and is defined as:
@@
where $\ell$ is the mixing length, $D_{\mathrm{MLT}}$ is the mixing diffusion coefficient, $\rho$ is the local density, $N_{A}$ is Avogadro’s number, $\langle\sigma v\rangle$ is the thermally averaged reaction rate, and $Y_{j}$ is the molar abundance of the interacting species. The diffusion coefficient is $D_{\mathrm{MLT}}=\frac{1}{3}v_{\mathrm{MLT}}\cdot\ell$, where $v_{\mathrm{MLT}}$ is the convective velocity and the mixing length is $\ell=\alpha\cdot H_{p}$ where $H_{p}$ is the pressure scale height and $\alpha$ is a free parameter (Joyce & Tayar, 2023).
@@
Existing massive star models that calculate $p$-nuclei nucleosynthesis are 1D and rely on mixing length theory (MLT) to describe convection. However, multi-dimensional hydrodynamic simulations of convective $\mathrm{O}$-shell burning predict higher convective velocities than MLT and show a gradual downturn in the mixing efficiency profile at shell boundaries (Meakin & Arnett, 2007; Jones et al., 2017). 3D simulations reveal features absent in 1D, such as asymmetric nuclear burning (Bazan & Arnett, 1994; Yadav et al., 2020), large-scale non-radial density asymmetries, and potentially lower $\mathrm{C}$-shell ingestion rates during $\mathrm{O}$-$\mathrm{Ne}$-$\mathrm{C}$ shell mergers (Andrassy et al., 2020; Yadav et al., 2020). In 1D models like those from Ritter et al. (2018b), $\mathrm{O}$-$\mathrm{C}$ shell mergers occur because the upper boundary of the $\mathrm{O}$-shell entrains $\mathrm{{}^{12}C}$ and $\mathrm{{}^{20}Ne}$ as it burns, flattening the entropy gradient. In contrast, 3D simulations by Rizzuti et al. (2024) find the lower boundary of the $\mathrm{C}$-shell extending downward and engulfing the $\mathrm{O}$-shell. Similarly, Yadav et al. (2020) show that entropy generation in nuclear burning hotspots within the $\mathrm{Ne}$-shell leads to downdrafts that raise the entropy in the $\mathrm{O}$-shell. Dynamic behaviour shortly before the core collapse for these supernova progenitors late in their evolution are not captured in 1D (Arnett & Meakin, 2011; Müller, 2016; Yadav et al., 2020). While 3D hydrodynamic simulations may not be solving all relevant equations, such as a robust nuclear network, 1D models fundamentally fail to represent the non-radial mixing and spherically asymmetric instabilities during $\mathrm{O}$-$\mathrm{C}$ shell mergers (Meakin & Arnett, 2006; Andrassy et al., 2020; Yadav et al., 2020).
@@
This has consequences for $p$-nuclei production. Nuclei with $A>110$ are primarily synthesized during the merger, not during explosive burning, regardless of the peak CCSN energy (Roberti et al., 2023, 2024). To explore the impact of the macrophysical uncertainties in the $\mathrm{O}$-shell during a merger, we adopt a 3D hydrodynamic inspired set of modifed radial mixing profiles and ingestion rates to determine the impact of mixing on the $\gamma$ process. We will also explore how varying the nuclear reaction rates impact the nucleosynthesis of the $p$ nuclei as the $\mathrm{O}$-shell is a convective-reactive environment.
@@
Section 2 describes the post-processing of the Ritter et al. (2018b) model using 3D hydrodynamic-inspired mixing profiles and assesses the impact of varying nuclear reaction rates. Section 3.1 examines how the convective-reactive environment produces the $p$ nuclei and the role of $\mathrm{C}$-shell ingestion. Sections 3.23.4 explore how 3D hydrodynamic insights affect $p$ nuclei production: 3.2 analyzes the impact of downturns and boosted velocities in the $\mathrm{O}$-shell, 3.3 investigates reduced $\mathrm{C}$-shell ingestion, and 3.4 evaluates dips in the mixing efficiency due to quenching. Section 3.5 presents the sensitivity to nuclear rates, their correlation with $p$ nuclei, and their relation to mixing profiles and velocities. Finally, Section 4 summarizes our findings and their implications.
2 Methodology
2.1 Initial Model and Post-Processing Setup
@@
Figure 1: Kippenhahn diagram showing the merger of the convective O and C-burning shells. The O-burning shell extends from $1.55\;\mathrm{\mathrm{M}_{\odot}}$ to $1.95\;\mathrm{\mathrm{M}_{\odot}}$, and the first ingested C-burning shell from $1.96\;\mathrm{\mathrm{M}_{\odot}}$ to $2.11\;\mathrm{\mathrm{M}_{\odot}}$. A red guideline has been provided to mark the thin radiative layer separating the O- and C-shell. Other convective regions are also C-burning shells. The merger onsets at $\log_{10}(t-t_{\mathrm{end}})/\mathrm{yr}\approx-3.85$ and reaches full extent at $\approx-4$. A black triangle marks where the initial composition is taken from and a white star marks the location where the ingested $\mathrm{C}$-shell material is taken from for this study.
@@
The detailed nucleosynthesis is calculated with the 1D multi-zone post-processing code mppnp (Pignatari et al., 2016). mppnp is a multi-zone post-processing code that uses stellar structure calculated by stellar evolution codes to calculate the full nucleosynthesis of a stellar model. mppnp treats mixing, nuclear burning, and ingestion separately rather than the coupled treatment by a code like MESA (Paxton et al., 2010). A convergence test was performed by decreasing the timesteps and increasing the number of mass zones. It found that a timestep of $\Delta t=0.01\;\mathrm{\mathrm{s}}$ and $400$ equidistant mass zones with $4$ additional zones at the bottom of the O-shell were sufficient to resolve both the burning and mixing timescales with a decreasing mixing efficiency profile. Calculations initially used a $5234$ isotope network, but many $n$-rich species were unnecessary. A network of only the necessary $1470$ isotopes was adopted, focusing on the $n$-deficient isotopes, for faster calculations. Isomeric states were not included in this network so $\mathrm{{}^{180m}Ta}$ is not calculated. A single simulation costs approximately $8$ hours on $40$ cores, for a total of $274$ core years for this work.
@@
The mppnp code calculates both the undecayed mass fraction as a function of mass and the mass-averaged contribubtions from radiogenic decay at a temperature of $T=100~{}\mathrm{MK}$ without explosive contributions. The mass fraction $X_{i}$ of species $i$ is defined as the ratio of the mass of that species to the total mass of the stellar material, such that $\sum_{i}X_{i}=1$ for all isotopes.
@@
To analyze the impact of macrophysical uncertainties and varying the nuclear reaction rates during the merger, isotopic mass fractions are taken just before the onset of merger at $\log_{10}(t-t_{\mathrm{end}})/\mathrm{yr}=-3.845$ from $m=1.8\;\mathrm{\mathrm{M}_{\odot}}$, and the ingested $\mathrm{C}$-shell material is taken from $m=2.0\;\mathrm{\mathrm{M}_{\odot}}$ at the same time, as shown in Figures 1 and 2. Earlier in the model, there is some $p$-nuclei production in the first convective $\mathrm{O}$-shell, which extends from $1.55\;\mathrm{\mathrm{M}_{\odot}}$ to $1.95\;\mathrm{\mathrm{M}_{\odot}}$ during $\log_{10}(t-t_{\mathrm{end}})/\mathrm{yr}=-1.76$ to $-2.16$. These nuclei are not processed by any further burning before the merger.
@@
where $X_{f}$ is the final mass-averaged decayed mass fraction of a species in the O-shell and $X_{i}$ is the inital mass fraction. The average overproduction factor is calculated as the arithmetic mean of $\mathrm{OP}$:
@@
where $N$ is the number of $p$ nuclei.
@@
MLT predictions of the radial mixing efficiency profile deviate from the more realistic predictions from 3D simulations. 3D convective $\mathrm{O}$-burning simulations show that the radial convective velocity profile gradually decreases near the shell boundary, in contrast to MLT predictions of a stiff boundary (Meakin & Arnett, 2007; Jones et al., 2017). This downturn is seen at both the bottom and top of convective shells (Herwig et al., 2006; Meakin & Arnett, 2007; Jones et al., 2017). The decrease occurs because mixing is driven by convective plumes in these simulations, rather than the idealized convective blobs in MLT. Plumes exhibit strong radial velocities in the interior of the convective region but lose their radial component as they reach the boundary, while non-radial velocity components increase. This behavior contrasts with MLT, which predicts a sharp drop to zero velocity at the boundary. Using Equation 4 from Jones et al. (2017), the downturn to the mixing efficiency profile can be implemented in 1D:
[…]
@@
where $\ell$ is the mixing length, $r_{0}$ is the Schwarzschild boundary at the bottom of the $\mathrm{O}$-shell, and an additional factor of $1/3$ is applied to match the MLT diffusion coefficient at the top of the shell.
@@
MLT also underpredicts the strength of the convective velocities in the $\mathrm{O}$-shell. Jones et al. (2017) found that convective velocities are stronger by a factor of $\sim 30$ compared to MLT. Andrassy et al. (2020) in their 3D C-shell entrainment simulations that their velocities could be up to $\sim 5$ times stronger than Jones et al. (2017) depending on the luminosity of $\mathrm{C}$- and $\mathrm{O}$-burning. It is possible the velocities could be even higher as these simulations do not include feedback from a nuclear network or treatment of radiation pressure (Jones et al., 2017; Andrassy et al., 2020). Rizzuti et al. (2024) finds that velocities are boosted by a factor of $\sim 10$ due to the feedback from new reactions with the ingested material in their 3D $\mathrm{O}$-$\mathrm{C}$ shell mergers. We implement this by applying a boost factor of 3×3$\times$3 ×, 10×10$\times$10 ×, and 50×50$\times$50 × to the convective velocities as shown in Figure 4.
@@
Figure 4: The diffusion coefficient profiles for the MLT and 3D-inspired gradual downturn scenarios. The dashed orange line is $D_{\mathrm{MLT}}$ and the dashed light grey line, dotted light blue line, solid grey line, and dashed dark blue line are the downturn profiles with boost factors of $1$, $3$, $10$, and $50$ respectively.
@@
The entrainment rate of $\mathrm{C}$-shell material could be lower during the merger depending on the strength of burning (Jones et al., 2017; Andrassy et al., 2020). To investigate the impact of this, a range of rates are considered: $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and a scenario with no entrainment. The maximum mass of the convective $\mathrm{C}$-shell in our model is $0.8\;\mathrm{\mathrm{M}_{\odot}}$, so for $110\;\mathrm{\mathrm{s}}$ of merging the maximum possible entrainment rate would be $7\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$ similar to Ritter et al. (2018a).
@@
1D and 3D simulations show that the convective profile can be quenched as material is ingested as entropy changes due to entrainment and burning (Iben, 1975; Sackmann et al., 1974; Herwig et al., 1999; Miller Bertolami et al., 2006; Herwig et al., 2011, 2014). As an example, during $\mathrm{H}$-ingestion into a $\mathrm{He}$-shell, the energy feedback from the ingested material could cause a split in the convective profile with very small amount of entrainment (Herwig et al., 2011). Herwig et al. (2014) found this effect could decrease the radial velocity profile and reduce the entrainment of species and labelled the event Global Oscillation of Shell Hydrogen-Ingestion (GOSH). Similar effects during C-shell entrainment could be possible, as Andrassy et al. (2020) found that strong oscillatory modes like GOSHs were present in their 3D simulations. Energy feedback events like this could explain supernova observations (Smith & Arnett, 2014). There is no clear prescription for how to implement this effect into 1D models, but it is clear that there would be decreased mixing because of a split. To investigate a possible convective quenching, we consider a GOSH-like event and a partial merger, where a Gaussian dip (but not full split) occurs in the MLT diffusion profile:

### v2

@@ Abstract
[…]
journal: ApJsoftware: The data and analysis tools are available on Zenodo under an open-source Creative Commons Attribution license: https://doi.org/10.5281/zenodo.17576026 (catalog doi:10.5281/zenodo.17576026). They are also available for download from the NuGrid collaboration website: download.nugridstars.org.
show]joshuaissa@uvic.ca
I Introduction
The origin of the solar pattern of the 35 stable $p$ nuclei is a long-standing problem for understanding nucleosynthesis. Burbidge et al. (1957) first identified these isotopes as a distinct group, produced primarily not by the $s$ or $r$ process, but instead by a $p$ process through $(p,\gamma)$, $(\gamma,p)$, and $(\gamma,n)$ reactions occurring during Type II supernovae and possibly Type I supernovae. Based on the first generations of stellar computational models, Arnould (1976) found that the $p$ process could be driven by all photo-disintegration reactions $(\gamma,n)$, $(\gamma,\alpha)$ and $(\gamma,p)$ during the most advanced evolutionary stages of massive stars. Woosley & Howard (1978) found these photo-disintegrations could create the distribution of $p$-process nuclei during the passage of the supernova shock over the internal progenitor structure, which they called the $\gamma$ process. Following works better defined $\gamma$ process production in Type II supernovae and more in general in core-collapse supernovae (CCSNe e.g., Prantzos et al., 1990; Rayet et al., 1995; Travaglio et al., 2018; Choplin et al., 2022; Roberti et al., 2023, 2024).
@@ I Introduction
The $\gamma$ process describes the flow of $(\gamma,n)$, $(\gamma,p)$, and $(\gamma,\alpha)$ reactions on the stable isotopic seeds that are already present in the shell at temperatures of $1.5\leq T\leq 3\times 10^{9}\;\mathrm{K}$ (Rauscher et al., 2013). However, the $\gamma$ process in massive stars underproduces not only the $\mathrm{Mo}$ and $\mathrm{Ru}$, but all $p$ nuclei with A=90130A=90{-}130 (Arnould & Goriely, 2003; Woosley & Heger, 2007). Travaglio et al. (2011) showed that the $\gamma$ process could produce all $p$ nuclei during a Type Ia supernova from the $s$-process material synthesized during stellar evolution without the underproduction of $\mathrm{{}^{92,94}Mo}$ and $\mathrm{{}^{96,98}Ru}$. Furthermore, Travaglio et al. (2015) found that modifying the distribution of $s$-process material significantly influenced the production of the $p$ nuclei, especially the heaviest ones, and that the lightest three were strongly dependent on the metallicity. Battino et al. (2020) additionally found that the H-flashes of rapidly accreting white dwarfs which undergo the $i$ process could modify the seed distribution to produce $p$ nuclei with $96<A<196$ by the $\gamma$ process during the subsequent SNIa.
Since the first isotopic classification made by Burbidge et al. (1957), it has been found that not all $p$ nuclei in the solar pattern are produced by a single process. Bisterzo et al. (2011) state that $\mathrm{{}^{152}Gd}$, $\mathrm{{}^{164}Er}$, and $\mathrm{{}^{180}Ta}$ have significant contributions of $70.5\%$, $75.5\%$, and $74.5\%$ from the $s$ process. Dillmann et al. (2008) found that $\mathrm{{}^{113}In}$ and $\mathrm{{}^{115}Sn}$ are made by $\beta$-decays after the $r$ process through isomeric states. Goriely et al. (2001) argue that $(\gamma,n)$ was too weak to produce $\mathrm{{}^{138}La}$, and instead that it is made by $\nu_{e}$-capture on $\mathrm{{}^{138}Ba}$ during the CCSN, and Arnould & Goriely (2003) similarly say that $\mathrm{{}^{180m}Ta}$ could also have $\nu$-induced contributions. Sieverding et al. (2018) also found that the $\nu$-process is important for the nucleosynthesis of $\mathrm{{}^{113}In}$, $\mathrm{{}^{138}La}$, and $\mathrm{{}^{180m}Ta}$.
The $\mathrm{O}$ shell where $p$ nuclei are produced during a merger is a convective-reactive environment where mixing and nuclear burning timescales are equal (Ritter et al., 2018a; Yadav et al., 2020). If there is significant energy released the flow can be modified (Dimotakis, 2005), such as $\mathrm{H}$-ingestion into $\mathrm{He}$-burning shell (Herwig et al., 2011, 2014) or $\mathrm{O}$-$\mathrm{C}$ shell mergers causing violent mixing (Andrassy et al., 2020; Yadav et al., 2020). The Damköhler number (Dimotakis, 2005) quantifies the ratio of these timescales and is defined as:
@@
where $\ell$ is the mixing length, $D_{\mathrm{MLT}}$ is the mixing diffusion coefficient, $\rho$ is the local density, $N_{A}$ is Avogadro’s number, $\langle\sigma v\rangle$ is the thermally averaged reaction rate, and $Y_{j}$ is the molar abundance of the interacting species. The diffusion coefficient is $D_{\mathrm{MLT}}=\frac{1}{3}v_{\mathrm{MLT}}\cdot\ell$, where $v_{\mathrm{MLT}}$ is the convective velocity and the mixing length is $\ell=\alpha\cdot H_{p}$ where $H_{p}$ is the pressure scale height and $\alpha$ is a free parameter (Vitense, 1953; Böhm-Vitense, 1958; Kippenhahn et al., 2013, as reviewed in Arnett et al., 2015).
@@
Existing massive star models that calculate $p$-nuclei nucleosynthesis are 1D and rely on mixing length theory (MLT) to describe convection. However, multi-dimensional hydrodynamic simulations of convective $\mathrm{O}$-shell burning predict higher convective velocities than MLT and show a gradual downturn in the mixing efficiency profile at shell boundaries (Meakin & Arnett, 2007; Jones et al., 2017). 3D simulations reveal features absent in 1D, such as asymmetric nuclear burning (Bazan & Arnett, 1994; Yadav et al., 2020), large-scale non-radial density asymmetries, and potentially lower $\mathrm{C}$-shell ingestion rates during $\mathrm{O}$-$\mathrm{Ne}$-$\mathrm{C}$ shell mergers (Andrassy et al., 2020; Yadav et al., 2020). In 1D models like those from Ritter et al. (2018b), $\mathrm{O}$-$\mathrm{C}$ shell mergers occur because the upper boundary of the $\mathrm{O}$ shell entrains $\mathrm{{}^{12}C}$ and $\mathrm{{}^{20}Ne}$ as it burns, flattening the entropy gradient. In contrast, 3D simulations by Rizzuti et al. (2024) find the lower boundary of the $\mathrm{C}$ shell extending downward and engulfing the $\mathrm{O}$ shell. Similarly, Yadav et al. (2020) show that entropy generation in nuclear burning hotspots within the $\mathrm{Ne}$ shell leads to downdrafts that raise the entropy in the $\mathrm{O}$ shell. Dynamic behaviour shortly before the core collapse for these supernova progenitors late in their evolution are not captured in 1D (Arnett & Meakin, 2011; Müller, 2016; Yadav et al., 2020). While 3D hydrodynamic simulations may not be solving all relevant equations, such as a robust nuclear network, 1D models fundamentally fail to represent the non-radial mixing and spherically asymmetric instabilities during $\mathrm{O}$-$\mathrm{C}$ shell mergers (Meakin & Arnett, 2006; Andrassy et al., 2020; Yadav et al., 2020).
@@
This has consequences for $p$-nuclei production. Nuclei with $A>110$ are primarily synthesized during the merger, not during explosive burning, regardless of the peak CCSN energy (Roberti et al., 2023, 2024). To explore the impact of the macrophysical uncertainties in the $\mathrm{O}$ shell during a merger, we adopt a 3D hydrodynamic inspired set of modified radial mixing profiles and ingestion rates to determine the impact of mixing on the $\gamma$ process. We will also explore how varying the nuclear reaction rates impact the nucleosynthesis of the $p$ nuclei as the $\mathrm{O}$ shell is a convective-reactive environment.
@@
Section II describes the post-processing of the Ritter et al. (2018b) model using 3D hydrodynamic-inspired mixing profiles and assesses the impact of varying nuclear reaction rates. Section III.1 examines how the convective-reactive environment produces the $p$ nuclei and the role of $\mathrm{C}$-shell ingestion. Sections III.2III.4 explore how 3D hydrodynamic insights affect $p$ nuclei production: III.2 analyzes the impact of downturns and boosted velocities in the $\mathrm{O}$ shell, III.3 investigates reduced $\mathrm{C}$-shell ingestion, and III.4 evaluates dips in the mixing efficiency due to quenching. Section III.5 presents the sensitivity to nuclear rates, their correlation with $p$ nuclei, and their relation to mixing profiles and velocities. Finally, Section IV summarizes our findings and their implications.
II Methodology
II.1 Initial Model and Post-Processing Setup
@@
Figure 1: Kippenhahn diagram showing the merger of the convective O and $\mathrm{C}$burning shells. The $\mathrm{O}$burning shell extends from $1.55\;\mathrm{\mathrm{M}_{\odot}}$ to $1.95\;\mathrm{\mathrm{M}_{\odot}}$, and the first ingested $\mathrm{C}$-burning shell from $1.96\;\mathrm{\mathrm{M}_{\odot}}$ to $2.11\;\mathrm{\mathrm{M}_{\odot}}$. A red guideline has been provided to mark the thin radiative layer separating the $\mathrm{O}$ and $\mathrm{C}$ shell. Other convective regions are also $\mathrm{C}$burning shells. The merger onsets at $\log_{10}(t-t_{\mathrm{end}})/\mathrm{yr}\approx-3.85$ and reaches full extent at $\approx-4$. A black triangle marks where the initial composition is taken from and a white star marks the location where the ingested $\mathrm{C}$-shell material is taken from for this study.
@@
The detailed nucleosynthesis is calculated with the 1D multi-zone post-processing code mppnp (Pignatari et al., 2016). mppnp is a multi-zone post-processing code that uses stellar structure calculated by stellar evolution codes to calculate the full nucleosynthesis of a stellar model. mppnp treats mixing, nuclear burning, and ingestion separately rather than the coupled treatment by a code like MESA (Paxton et al., 2010). A convergence test was performed by decreasing the timesteps and increasing the number of mass zones. It found that a timestep of $\Delta t=0.01\;\mathrm{\mathrm{s}}$ and $400$ equidistant mass zones with $4$ additional zones at the bottom of the $\mathrm{O}$ shell were sufficient to resolve both the burning and mixing timescales with a decreasing mixing efficiency profile. Calculations initially used a $5234$ isotope network, but many $n$-rich species were unnecessary. A network of only the necessary $1470$ isotopes was adopted, focusing on the $n$-deficient isotopes, for faster calculations. Isomeric states were not included in this network so $\mathrm{{}^{180m}Ta}$ is not calculated. Entrainment of $\mathrm{C}$-shell material is treated the same as described in Denissenkov et al. (2019) and does not depend on the time step. A single simulation costs approximately $8$ hours on $40$ cores, for a total of $274$ core years for this work.
@@
The mppnp code calculates both the undecayed mass fraction as a function of mass and the decayed mass-averaged mass fraction at a temperature of $T=100~\mathrm{MK}$ without explosive contributions. The mass fraction $X_{i}$ of species $i$ is defined as the ratio of the mass of that species to the total mass of the stellar material, such that $\sum_{i}X_{i}=1$ for all isotopes.
@@
To analyze the impact of macrophysical uncertainties and varying the nuclear reaction rates during the merger, isotopic mass fractions are taken just before the onset of merger at $\log_{10}(t-t_{\mathrm{end}})/\mathrm{yr}=-3.845$ from $m=1.8\;\mathrm{\mathrm{M}_{\odot}}$, and the ingested $\mathrm{C}$-shell material is taken from $m=2.0\;\mathrm{\mathrm{M}_{\odot}}$ at the same time, as shown in s 1 and 2. Earlier in the model, there is some $p$-nuclei production in the first convective $\mathrm{O}$ shell, which extends from $1.55\;\mathrm{\mathrm{M}_{\odot}}$ to $1.95\;\mathrm{\mathrm{M}_{\odot}}$ during $\log_{10}(t-t_{\mathrm{end}})/\mathrm{yr}=-1.76$ to 2.16-2.16. These nuclei are not processed by any further burning before the merger.
@@
where $X_{f}$ is the final mass-averaged decayed mass fraction of a species in the $\mathrm{O}$ shell and $X_{i}$ is the inital mass fraction. The average overproduction factor is calculated as the arithmetic mean of $\mathrm{OP}$:
@@
where $N$ is the number of $p$ nuclei. The mean logarithmic value is defined similar to hs and ls for the $s$ process as done by Busso et al. (1999).
@@
MLT predictions of the radial mixing efficiency profile deviate from the more realistic predictions from 3D simulations. 3D convective $\mathrm{O}$ burning simulations show that the radial convective velocity profile gradually decreases near the shell boundary, in contrast to MLT predictions of a stiff boundary (Meakin & Arnett, 2007; Jones et al., 2017). This downturn is seen at both the bottom and top of convective shells (Herwig et al., 2006; Meakin & Arnett, 2007; Jones et al., 2017). The decrease occurs because mixing is driven by convective plumes in these simulations, rather than the idealized convective blobs in MLT. Plumes exhibit strong radial velocities in the interior of the convective region but lose their radial component as they reach the boundary, while non-radial velocity components increase. This behavior contrasts with MLT, which predicts a sharp drop to zero velocity at the boundary. Using Equation 4 from Jones et al. (2017), the downturn to the mixing efficiency profile can be implemented in 1D:
$D_{\mathrm{3D\text{-}insp.}}=\frac{1}{3}v_{\mathrm{MLT}}\times\min{(\ell,r-r_{0})}$ (5)
@@
where $\ell$ is the mixing length, $r_{0}$ is the Schwarzschild boundary at the bottom of the $\mathrm{O}$ shell, and an additional factor of $1/3$ is applied to match the MLT diffusion coefficient at the top of the shell.
@@
MLT also underpredicts the strength of the convective velocities in the $\mathrm{O}$ shell. Jones et al. (2017) found that convective velocities are stronger by a factor of $\sim 30$ compared to MLT. Andrassy et al. (2020) in their 3D $\mathrm{C}$-shell entrainment simulations show that their velocities could be up to $\sim 5$ times stronger than Jones et al. (2017) depending on the luminosity of $\mathrm{C}$ and $\mathrm{O}$ burning. It is possible the velocities could be even higher as these simulations do not include feedback from a nuclear network or treatment of radiation pressure (Jones et al., 2017; Andrassy et al., 2020). Rizzuti et al. (2024) finds that velocities are boosted by a factor of $\sim 10$ due to the feedback from new reactions with the ingested material in their 3D $\mathrm{O}$-$\mathrm{C}$ shell mergers. We implement this by applying a boost factor of $3\times$, $10\times$, and $50\times$ to the convective velocities as shown in Figure 4.
@@
Figure 4: The diffusion coefficient profiles for the MLT and 3D-inspired gradual downturn scenarios. The dashed orange line is $D_{\mathrm{MLT}}$ and the dashed light grey line, dotted light blue line, solid grey line, and dashed dark blue line are the downturn profiles with boost factors of $1$, 33, $10$, and $50$ respectively.
@@
The entrainment rate of $\mathrm{C}$-shell material could be lower during the merger depending on the strength of burning (Jones et al., 2017; Andrassy et al., 2020). To investigate the impact of this, a range of rates are considered: $4\times 10^{-5}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-4}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, $4\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$, and a scenario with no entrainment. The maximum mass of the convective $\mathrm{C}$-shell in our model is $0.8\;\mathrm{\mathrm{M}_{\odot}}$, and since our simulation goes from $\log_{10}(t-t_{\mathrm{end}})/\mathrm{yr}=-3.856$ to -3.845 ($110\;\mathrm{\mathrm{s}}$), the maximum entrainment rate would be $7\times 10^{-3}\;\mathrm{\mathrm{M}_{\odot}\mathrm{s}^{-1}}$ similar to Ritter et al. (2018a).
@@
1D and 3D simulations show that the convective profile can be quenched as material is ingested as entropy changes due to entrainment and burning (Iben, 1975; Sackmann et al., 1974; Herwig et al., 1999; Miller Bertolami et al., 2006; Herwig et al., 2011, 2014). As an example, during $\mathrm{H}$-ingestion into a $\mathrm{He}$-shell, the energy feedback from the ingested protons quickly burning can cause a split in the convective profile with a very small amount of entrainment (Herwig et al., 2011). Herwig et al. (2014) found this effect could decrease the radial velocity profile and reduce the entrainment of species and labelled the event Global Oscillation of Shell Hydrogen-Ingestion (GOSH). Similar effects during $\mathrm{C}$-shell entrainment could be possible, as Andrassy et al. (2020) found that strong oscillatory modes like GOSHs were present in their 3D simulations. Energy feedback events like this could explain supernova observations (Smith & Arnett, 2014). There is no clear prescription for how to implement this effect into 1D models, but it is clear that there would be decreased mixing because of a split. To investigate a possible convective quenching, we consider a GOSH-like event and a partial merger, where a Gaussian dip (but not full split) occurs in the MLT diffusion profile: